Unbiased Expectations Theory One-year Treasury bills currently earn 5.35 percent. You expect that one year from now, one-year Treasury bill rates will increase to 5.50 percent. If the unbiased expectations theory is correct, what should the current rate be on two-year Treasury securities?

  • 5.5000%

  • 5.4250%

  • 10.8500%

  • 5.3500%
Answer #1

Based on Unbiased expectations theory,

(1 + 2 Yr rate)2 = (1 + 1 Yr rate) * (1 + 1 Yr Rate 1 Yr from now)

(1 + 2 Yr rate)2 = (1 + 5.35%) * (1 + 5.50%)

(1 + 2 Yr rate)2 = 1.111443

(1 + 2 Yr rate) = 1.054250

2 Yr rate = 0.054250

2 Yr rate = 5.4250%